# Beyond One-Hot Labels: Semantic Mixing for Model Calibration

### v1

@@ 5 Conclusion
In conclusion, CSM is a novel framework designed to bridge the optimization gap between the calibration objective and the data of full certainty. By generating augmented samples with semantic mixing and reannotating them with confidence scores via diffusion models, CSM offers a more accurate alignment between model predictions and their true likelihoods. Our exploration of balanced loss functions further enhances the new data representation paradigm. Theoretical and practical evidence validate the effectiveness of CSM for strong sample-label augmentations and its accurate learning objective.
@@ Appendix C More Experimental Results
Figure 10: Analysis of the amount of augmented data and hyperparameter $s\in[2.0,6.0]$. (a) OE and UE w.r.t. different $s$ values. (b) ECE/AECE and their post-temperature values w.r.t different $s$ values. (c) Classification ACC, ECE, and AECE on CIFAR-100 w.r.t different number of augmented samples relative to the dataset size.

### v2

@@ 5 Conclusion
In conclusion, CSM is a novel framework designed to bridge the optimization gap between the calibration objective and the data of full certainty. By generating augmented samples with semantic mixing and reannotating them with confidence scores via diffusion models, CSM enables a more accurate alignment between model predictions and their true likelihoods. Our exploration of balanced loss functions further enhances the new data representation paradigm, enhancing it an integrated pipeline for superior model calibration. Theoretical and practical evidence validate the effectiveness of CSM for strong sample-label augmentations. Meanwhile, various ablation results demonstrate CSM’s balanced learning of true confidence levels. The framework’s ability to acquire meaningful augmentation positions it as a novel baseline for semantic-aware confidence calibration.
Acknowledgements
This work was supported in part by the Start-up Grant (No. 9610680) of the City University of Hong Kong, Young Scientist Fund (No. 62406265) of NSFC, and the Australian Research Council under Projects DP240101848 and FT230100549.
@@ Appendix C More Experimental Results
Figure 8: Analysis of the amount of augmented data: Classification ACC, ECE, and AECE on CIFAR-100 w.r.t different number of augmented samples relative to the dataset size.
Table 11: Calibration Comparison with Different Number of Augmented Samples per Training Sample.
$N_{\text{augs}}$ 1 2 3
ECE: CIFAR-10 0.83 0.54 0.39
ECE: CIFAR-100 2.07 1.29 1.74
Table 12: Comparison of the Estimated Training Time.
Method CE MbLS Mixup RegMixup RankMixup Ours Ours(EQ-DATA)
Training Time 2.63h 2.65h 3.48h 3.50h 4.30h 4.28h 2.64h
@@
Number of Augmented Samples per Training Sample We provide more analyses about the number of augmented samples per training sample (denoted as $N_{aug}$) with results in Table 11. It can be observed that adding the number of accompanied augmentations per dataset sample can generally improve the final calibration performance. This is because a larger number of $N_{augs}$ can sample more sufficient proximal data for training, better filling the domain space and providing more accurate confidence estimation. However, simply using larger $N_{augs}$s could also raise the computational overhead and slow down the training.
Estimated Training Time We provide a comparison of the estimated training time in Table 12. One can see the number of augmented samples per batch is the main factor influencing the training time. CSM outperforms others in calibration while maintaining reasonable speed. Even with equalized training samples, i.e., EQ-DATA, it achieves competitive calibration performance. Augmented samples need no re-generation across model, objective, or annotation changes, enabling efficient modular study. We run CSM with a single RTX A4000 device.